OpenStudy (anonymous):

Simplify the given expression. (6 + 2i) − (8 − 3i)

OpenStudy (usukidoll):

ahhh must be in ai+b or a+bi form

OpenStudy (anonymous):

−2 + 5i −2 − i 8 + 5i 8 − 5i

OpenStudy (usukidoll):

distribute the negative on the right hand side of the equation ... combine like terms.. there's the format ^^

OpenStudy (usukidoll):

(6 + 2i) − (8 − 3i) 6+2i-8+3i = -2+5i

OpenStudy (anonymous):

how'd you add them together? i tried foil but i didnt get the right thing

OpenStudy (usukidoll):

no there's no foil in here. just normal subtraction and combining like terms

OpenStudy (usukidoll):

foil works for (x+y)(x-y) no sign in the middle

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

can you help with one more? Simplify the given expression. 2 divided by the quantity of 2 plus 5i

OpenStudy (usukidoll):

hmm :/ I can't see it very well

OpenStudy (usukidoll):

argh I forgot how those work...

OpenStudy (usukidoll):

I know you can't split them up... nooo that breaks the rules

OpenStudy (mathmale):

Couple of things to know here: 1) i=Sqrt(-1), i^2=-1 2) (a+b)(a-b) = a^2 - b^2, with no middle term 3) the goal of this problem is to eliminate the complex number 2 +5i from the denominator. 4) 2-5i is called the "conjugate" of 2+5i. 5) Applying principles from (1) and (2), above, (2-5i)(2+5i)=4-25i^2 which simplifies to 4-25(-1) = 29. 6) multiply both numerator and denominator of the given expression by the conjugate in (4) and simplify the result. 2(2-5i) 2(2-5i) ------- = --------- is the simplest form the answer you want. 4-25(-1) 29 There's no real benefit from multiplying out 2(2-5i).

OpenStudy (mathmale):

Notice how the denominator is now real, whereas before it was complex.

OpenStudy (anonymous):

so you just multiply the numerator by the conjugate? then simplify fro mthere?

OpenStudy (mathmale):

Also notice that the first problem you presented was one in subtraction of complex numbers; the second was more complicated because of the multiplication of complex numbers required. @skipper16: Multiply BOTH the numerator and the denominator of the given fraction by the conjugate (2-5i).

OpenStudy (mathmale):

This process may be referred to as "rationalizing the denominator".

OpenStudy (usukidoll):

now I remember...

OpenStudy (anonymous):

Hmm

OpenStudy (mathmale):

Thank you very much for the medal.

OpenStudy (anonymous):

you're welcome thanks for the help

OpenStudy (usukidoll):

:O where's my medal?